OpenStudy (vuriffy):

(9-3 x 4 / 6 + 3)/((8/4)^2 + 3 x 7 - 20) Enjoy. Just a little logic question for these lovely math people. Trust me, it's easy.

Nnesha (nnesha):

PEMDAS

OpenStudy (vuriffy):

Oh I know, hehe. I already did it in my head, lol. I wanted someone to have fun with it.

Nnesha (nnesha):

ik but it would be great if you take a screenshot or rewrite n Latex use the VV equation button :=))

OpenStudy (vuriffy):

\[(9-3 \times 4 \div 6 + 3) / ((8\div4)^2 + 3 \times 7 -20)\]

OpenStudy (vuriffy):

Do you mean this?

Nnesha (nnesha):

better :=)) i don't really trust people but i trust you it is EASY :d lel

Nnesha (nnesha):

\[\frac{ 9 \color{REd}{-3 \times 4 }\div 6 +3}{ (8 \div 4)^2 + 3 \times 7 -20) }\] according to the PEMDAS first Parentheses exponent multiply divide add subtract there is no parentheses or exponent at the numerator so first multiply \[\frac{ 9 \color{REd}{-12 }\div 6 +3}{ (8 \div 4)^2 + 3 \times 7 -20) }\] 2nd divide \[\frac{ 9 \color{blue}{-12 \div 6} +3}{ (8 \div 4)^2 + 3 \times 7 -20) }\] \[\frac{ 9 \color{blue}{-2} +3}{ (8 \div 4)^2 + 3 \times 7 -20) }\] and then simple add& subtract

OpenStudy (zzr0ck3r):

lol at this being an exercise in logic. We could make it about logic and explain every step and why we can use these steps, and which axioms are we 'abusing' to use these steps. :)

OpenStudy (vuriffy):

@zzr0ck3r Your response made my day that day. Haha.